Use sum or difference identities to find the exact value of each trigonometric function.
4.csc(915)degrees
6.sec(15)degrees
I tried this one on my own, could someone please, if possible verify that it is a right answer. sin (-165)degrees = (-radical6 + radical 2)/(-4)
Thanks for helping me!

Answer by stanbon(48558)

Use sum or difference identities to find the exact value of each trigonometric function.
4.csc(915)degrees
915-2*360 = 195
--------
csc(915) = csc(105+90) = [sin(105+90)]^-1 = [sin105*cos90 + cos105*sin90]^-1
[0 + cos105]^-1 = [-0.258819...]^-1 = -3.8637...
----------------
6.sec(15)degrees = [cos(45-30)]^-1
= [cos45cos30+sin45sin30]^-1
= [(sqrt(2)/2)(sqrt3)/2) + (sqrt2/2)(1/2)]^-1
= [sqrt(6)/4 + sqrt(2)/4]^-1
= [(sqrt(6)+ sqrt(2)/4]^-1
= 1.0353...
================
Cheers,
Stan H.